40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Message ID: 1870     Entry time: Tue Dec 1 18:37:09 2020
Author: Paco 
Type: DailyProgress 
Category: Optics 
Subject: OPO pump steering 

Enter lab ~09:20. Today I spent a while looking at the broadband EOM drivers used in CTN (presently optimized for 37 MHz) and installed the preceding steering and power control (half waveplate + pbs) optics. The beam path for the OPO pump beam is now set to 3 inches (note the NPRO head is nominally 4 inch above the table).

ELOG V3.1.3-